A cylindrical spring is found to elongate by 3 cm when stretched by a force of 500N.


Determine: 

 a) its force constant

 b) the work done by a 600N force in elongating the spring


1
Expert's answer
2021-03-23T11:11:30-0400

a) "F=k\Delta x\to k=F/\Delta x=500/0.03=16667\ (N/m)"


b) "W=PE=\frac{k(\Delta x)^2}{2}=\frac{k(F/k)^2}{2}=\frac{16667\cdot(600/16667)^2}{2}=10.8\ (J)"